a boy moves 5m east and then 5m north.Its displacement to distance ratio is?

Grade:12th pass

1 Answers

Khimraj
5 years ago
displacement will be 5\sqrt{2} m
distance = 5+5  = 10
so displacement/distance       = 1/\sqrt{2}
Hope it clears................................